OpenStudy (anonymous):

Find all real solutions of (x^2 + 7)^3/2 = 64

OpenStudy (anonymous):

taking square both sides equation becomes. \[(x^{2}+7)^{3} = 64*64\] now taking cube root \[(x^{2}+7) = 4*4=16\] \[x^{2}=9 \implies x=+3,-3\] so solution of above equation are 3 and -3